. Calculate the concentration of the following solutions:

     a. mass percent (m/m) for the solute: 75 g of Na2SO4 and 175 g of H2O

     b. mass/volume percent (m/v) for the solute: 54 g of KCl in 325 mL of KCl solution
Answer #1

A) Total mass of solution = 75 + 175 = 250 g

(m/m) % concentration = 75*100/250 = 30 % ....Answer

B)

(m/V) % concentration = 54 * 100 / 325 = 16.62 % ....Answer